Gas Law problems

1) What is the volume of 2.75 moles of phosgene gas (COCl2) at 298.15 K and 1.25 atm?

2) Assuming ideal behavior, what is the volume of 2.00 g of O2 at STP?

3) Calculate the density of fluorine gas at 22.0 ºC and 103.2 kPa.

4) Given the same set of conditions, how much longer will it take for SO3 to effuse the same distance as NO2?
Answer #1

1)

Given:

P = 1.25 atm

n = 2.75 mol

T = 298.15 K

use:

P * V = n*R*T

1.25 atm * V = 2.75 mol* 0.08206 atm.L/mol.K * 298.15 K

V = 53.8256 L

Answer: 53.8 L

2)

Molar mass of O2 = 32 g/mol

mass(O2)= 2 g

use:

number of mol of O2,

n = mass of O2/molar mass of O2

=(2 g)/(32 g/mol)

= 6.25*10^-2 mol

Given:

P = 1.0 atm

n = 0.0625 mol

T = 273.0 K

use:

P * V = n*R*T

1 atm * V = 0.0625 mol* 0.08206 atm.L/mol.K * 273 K

V = 1.4001 L

Answer: 1.40 L

Only 1 question at a time please